Vocaloid:

need some help deriving an equation for the mass of an ion in a mass spectrometer in terms of voltage, magnetic field strength, radius of particle deflection, and charge

Vocaloid:

I am given: qvB = mv^2/r qV = (1/2)mv^2 the sol'n is m = \(\frac{ qB^2*r^2 }{ 2V }\) I feel like I'm missing something obvious but I can't quite get that solution

nuts:

\(qvB = \frac{mv^2}{r}\) \(qV = (1/2)mv^2\) \(v^2=\frac{2qV}{m};\) \(qvB=\frac{m(2qV/m)}{r}=2qV/r\) \(vB=\frac{2V}{r}\) \(v=\frac{2V}{rB}\) \(qV=(1/2)mv^2\) \(\frac{2qV}{v^2}=m\) \(m=\frac{2qV}{(\frac{2V}{rB})^2}=\frac{2qVr^2B^2}{4V^2}=\frac{qB^2r^2}{2V}\) \(\text{as desired}\) now please unblock me so i can send you more annoying spam so you can kick me from all subjects repeatedly and block me again :D

nuts:

@Vocaloid

Vocaloid:

thank you

nuts:

Look I owe you an apology for being obnoxious. I will do better.